OpenStudy (anonymous):

do you know the formula for sec in geometry i am so confused. Ex. sec angle B = 3/7

OpenStudy (anonymous):

secX=1/cosX

OpenStudy (anonymous):

so 1/cosx =3/7

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

can you teach me how to do this?

OpenStudy (anonymous):

Given the triangle below, which of the following is a correct statement?

OpenStudy (anonymous):

cotC=1/2 is true

OpenStudy (anonymous):

because tanC=2 hence since tan = 1/cot then cot =1/tan = 1/2

OpenStudy (anonymous):

ohhh

OpenStudy (anonymous):

I was so confused but nit anymore thank you very much

OpenStudy (anonymous):

no worries. i gotta go to bed now mate. good luck!